The square of a number added to the original number is 72.
Determine all possible values of the number.

Answer by checkley71(8403)

X+X^2=72
X^2+X-72=0
(X+9)(X-8)=0
X-8=0
X=8 ANSWER.
PROOF
8+8*8=72
8+64=72
72=72